# Projected Augmented Waves (PAW) inspired mixed basis sets for small molecules

### v3

@@ Abstract
The success behind many pseudopotential methods such as Projected Augmented Waves (PAW) and Phillips-Kleinman pseudopotential methods is that these methods are all electron methods in disguise. For the Phillips Kleinman and PAW methods we show that there is an all electron formulation. In the all electron formulation there are regular low wavevector basis wave functions and several, specially chosen, high wavevector basis wave functions that are specialized to the atomic environment. Using this as inspiration here we propose a new, PAW method inspired, basis set for small molecules, where we use the recently introduced LO or lo (Localized orbitals) basis set is paired with a Gaussian basis, Gaussian type orbitals (GTO), for a hybrid basis for small molecules. Contracted Gaussian functions are also possible (CGFs). We also show how to extend this idea to Slater basis set, Slater type orbitals (STO), instead of GTO orbitals combined with LO or lo orbitals.
@@ I Introduction
It is of paramount importance to make further progress on the basis set problem for Density Functional Theory (DFT) for small molecules. Small molecules larger then diatomic: carbon dioxide, water, ammonia, methane, ethyne, ethene, ethane to name a few have a variety of uses in industry and engineering [2, 1]. There has been as much progress in optimizing basis sets for small molecule DFT calculations, similarly to basis sets that have paved breakthroughs in solid state DFT [3, 4, 5, 9, 11, 12, 17, 14, 13, 8, 7, 6, 1, 10, 15, 16]. Much progress has been made using ideas based on Linear Combinations of Atomic Orbitals (LCAO) [19, 2, 18] where a good basis for the entire molecule consists of linear combinations of basis functions for the individual atoms. There are many candidate LCAO basis sets:
@@
  • Slater basis sets where Slater Type Orbitals (STOs) are used in the LCAO method. These are spherical harmonics multiplied by exponentially decaying radial wave functions. It is possible to have multiple wave functions per angular momentum channel - so called multi-zeta basis sets. STO’s have advantages over GTOs as they have the correct exponentially decaying behavior at infinity and the right Kato cusp at the origin [18, 2], however four center integrals are much harder.
@@
Furthermore plane waves, outside the LCAO concept, are often used to study periodically arranged arrays of molecules, with the distance between molecules on the order of 10 $nm$ so that interactions between the molecules may be well neglected. The most common electronic structure method for this situation (plane waves) is Projected Augmented Waves (PAW) [24, 25, 3, 4].
Here we focus on the LCAO method. Inspired by pseudopotential methods in DFT we propose one such (LCAO) hybrid basis set. Pseudopotentials at their core map a problem into another problem. The exact all electron wavefunction for the system is mapped onto a smooth wavefunction that may well be represented by a small number of plane waves, that is it is on its face not all electron. Two of the pre-eminent examples of pseudopotentials are the Phillips Kleinman and Projected Augmented Waves (PAW) pseudopotentials. Here we show that these pseudopotential methods are all electron methods in disguise, where the all electron problem includes smooth plane wave wave functions and specialized wave functions for orbitals near atomic nuclei. Inspired by this observation, and the extensive numerical success of the PAW method [3, 4] we propose new basis sets for small molecules, where localized LO or lo [26, 5, 6, 1] basis wave functions are augmented with Gaussian wave functions - GTOs - (where Gaussians have been extensively used for the problem and various configuration integrals for them are well known [18, 2, 3]) or STO’s for enhanced accuracy. This is done in a manner similar to Phillips-Kleinman and PAW all electron basis sets. We propose that this new hybrid basis set is efficient for DFT and Hartree-Fock (HF) calculations of small molecules.
@@ II Many pseudopotential methods are all electron methods in disguise (motivation)
In this section we reformulate PAW and the Phillips-Kleinman method as all electron methods. We show there are plane waves and special wave functions that are chosen to represent states near the core. We show the two methods to be equivalent in their all electron formulation depending on the choice of the localized wave functions.
@@ II.1 Phillips-Kleinman as an all electron method in disguise
We now recall the Orthogonalized Plane Wave (OPW) wave functions used for the Phillips-Kleinman method. We know that:
@@
$-\sum_{i}\left\langle\phi_{i}\left(\mathbf{r}\right)\mid\frac{1}{\sqrt{V}}\exp\left(i\left(\mathbf{k}+\mathbf{K}\right)\cdot\mathbf{r}\right)\right\rangle\left|\phi_{i}\left(\mathbf{r}\right)\right\rangle$ (1)
Here $\left|\phi_{i}\left(\mathbf{r}\right)\right\rangle$ are localized wave functions relevant to the single site problem. Here $V$ is the volume of the primitive lattice cell and $\mathbf{k}$is a wavevector in the first Brillouin zone while $\mathbf{K}$ is a reciprocal lattice vector. Now the main claim of the Phillips-Kleinman method (greatly numerically supported) is that for a reasonable number of plane waves the exact Khon Sham (KS) wavefunction $\left|\psi_{n}\right\rangle$ ca be written as
$\left|\psi_{n}\right\rangle\cong\sum_{\mathbf{k},\mathbf{K}}c_{\mathbf{k}}^{\mathbf{K}}\left|\phi_{OPW}\left(\mathbf{k}+\mathbf{K}\right)\right\rangle$ (2)
for some reasonable cutoffs $\mathbf{K}_{\mathbf{max}}$ that means that
@@
$\subset Span\left\{\left\{\left|\frac{1}{\sqrt{V}}\exp\left(i\left(\mathbf{k}+\mathbf{K}\right)\cdot\mathbf{r}\right)\right\rangle\right\},\left\{\left|\phi_{i}\left(\mathbf{r}\right)\right\rangle\right\}\right\}$ (3)
@@
$\left\{\left\{\left|\frac{1}{\sqrt{V}}\exp\left(i\left(\mathbf{k}+\mathbf{K}\right)\cdot\mathbf{r}\right)\right\rangle\right\},\left\{\left|\phi_{i}\left(\mathbf{r}\right)\right\rangle\right\}\right\}$ (4)
As such Phillips-Kleinman is an all electron method in disguise. We shall see below that PAW is very similar.
@@ II.2.1 PAW review
We would like to study highly oscillatory wave functions near the nucleus of an atom using plane waves. In order to obtain a low cutoff of say ~50 Rydberg in energy and as such a tolerable basis set we must have that the wavefunction we represent using plane waves is very smooth. The key idea of PAW is then to introduce:
@@
where $\hat{\mathcal{T}}$ is some linear transformation (specified below) and $\left|\psi_{n}\right\rangle$ is the exact Khon Sham wave functions (solutions of the KS equations) while $\left|\tilde{\psi}_{n}\right\rangle$ are the wave functions we approximate efficiently using plane waves. Now suppose there is a set of augmentation spheres and we wish for $\hat{\mathcal{T}}$ to be the identity outside the augmentation spheres (centered around nuclei) , so we write:
@@
Here $\mu$ are the atomic augmentation spheres. Now we want the wavefunction $\left|\tilde{\psi}\right\rangle$ to be smooth inside the sphere and we know that the exact KS wavefunction is close enough to the single atom, atomic wavefunction $\left|\phi_{i}\right\rangle$. As such if we choose $S_{\mathbf{\mu}}$ to be
$S_{\mu}\left|\psi_{n}\right\rangle=\sum c_{i}^{n}\left|\phi_{i}\right\rangle-\sum c_{i}^{n}\left|\tilde{\phi}_{i}\right\rangle$ (7)
where $\left|\tilde{\phi}_{i}\right\rangle$ are smooth wave functions. Where we choose
$c_{i}^{n}=\left\langle\tilde{p}_{i}\mid\tilde{\psi}_{n}\right\rangle$ (8)
Here $\left|\tilde{p}_{i}\right\rangle$ are the projector wave functions. As such we have that:
$\left|\psi_{n}\right\rangle=\hat{\mathcal{T}}\left|\tilde{\psi}_{n}\right\rangle=\left|\tilde{\psi}_{n}\right\rangle+\sum_{i}\left(\left|\phi_{i}\right\rangle-\left|\tilde{\phi}_{i}\right\rangle\right)\left\langle\tilde{p}_{i}\mid\tilde{\psi}_{n}\right\rangle$ (9)
@@ II.2.2 Reformulation of PAW as an all electron method
$\left|\psi_{n}\right\rangle\cong\left|\tilde{\psi}_{n}\right\rangle+\sum_{i}\left(\left|\phi_{i}\right\rangle-\left|\tilde{\phi}_{i}\right\rangle\right)\left\langle\tilde{p}_{i}\mid\tilde{\psi}_{n}\right\rangle.$ (10)
@@
$\left|\psi_{n}\right\rangle\in Span\left\{\left\{\frac{1}{\sqrt{V}}\exp\left(i\left(\mathbf{k}+\mathbf{K}\right)\cdot\mathbf{r}\right)\right\},\left\{\left|\phi_{i}\right\rangle\right\},\left\{\left|\tilde{\phi}_{i}\right\rangle\right\}\right\}$ (11)
@@
$\left|\tilde{\phi}_{i}\right\rangle\in Span\left\{\frac{1}{\sqrt{V}}\exp\left(i\left(\mathbf{k}+\mathbf{K}\right)\cdot\mathbf{r}\right)\right\}$ (12)
@@
$\left|\psi_{n}\right\rangle\in Span\left\{\left\{\frac{1}{\sqrt{V}}\exp\left(i\left(\mathbf{k}+\mathbf{K}\right)\cdot\mathbf{r}\right)\right\},\left\{\left|\phi_{i}\left(\mathbf{r}\right)\right\rangle\right\}\right\}$ (13)
for a reasonable number of plane waves. Since the basis set of PAW and Phillips-Kleinman are similar both methods should yield in their all electron formulation similar accuracy and basis of the same size. They are equivalent within the all electron formulation when they use the same localized orbitals.
@@ III GLO, SLO and PLO basis sets
Here we would like to work backwards we wish to formulate an all electron method as the inverse of a pseudopotential method where we use specialized wave functions near the nuclei and many regular wave functions for the whole space. We focus on small molecules and crystalline solids. Below we will simply describe the relevant basis sets.
@@ III.1 Glo and GLO for small molecules
Here G in Glo and GLO stands for Gaussians while lo and LO for localized orbitals. We consider the lo basis set [26]:
@@
where $S_{l}^{\mu}$ is chosen to make the wavefunction continuous everywhere. Here:
@@
and $\bar{V}_{KS}\left(r\right)$ is the spherically average KS potential or HF single particle potential. Furthermore $E_{l}$ is chosen at the binding energy of the single atom system (other options are possible). Alternatively we can use LO basis sets [26]:
@@
where $B_{l}^{\mu}$ and $S_{l}^{\mu}$ are chosen to make the wavefunction continuous and continuously differentiable everywhere. Here $\dot{u}_{l}^{E_{l}}=\frac{\partial}{\partial E}u_{l}^{E_{l}}$. Higher derivative terms are also possible [26], also regular LO and lo is possible [5, 6, 1]. We also add a large number of Gaussians to our basis set:
@@
to the molecular basis set. $G_{l,m,\sigma}^{\mu}$, these help make configuration integrals cheap. Here $\mathcal{N}_{l}$ is a normalization constant and cartesian Gaussians are also possible. This is a basis set for many small molecules. Furthermore Gaussians Coulomb integrals may be efficiently computed [28, 18] so there are “few” coulomb integrals involving complicated functions [29]. In particular if we consider the number of LO or lo orbitals as order one this becomes a three center problem at worst. We note that contracted Gaussian Functions (CGF) are also possible [27, 18]

### v4

# Projected Augmented Waves (PAW) motivated mixed basis sets for small molecules

@@ Abstract
The success behind many pseudopotential methods, such as Projected Augmented Waves (PAW) and Phillips-Kleinman pseudopotential method, is that these methods are all electron methods in disguise. For the Phillips Kleinman and PAW pseudopotential methods we show that there is an explicit all electron reformulation. In the all electron reformulation, in the basis set, there are regular low wavevector basis wave functions and several, specially chosen, high wavevector basis wave functions that are specialized to the atomic environment relevant to nuclei of the substances studied. Using this as motivation, here we propose a new, PAW method motivated, basis set for small molecules, where we use the LO or lo (Localized Orbitals) basis wave functions, and pair them with a Gaussian basis, e.g. Gaussian Type Orbitals (GTO), for a hybrid basis for small molecules. Contracted Gaussian functions are also possible (CGFs). Several different, but related, LO (lo) basis wave functions are considered. We also show how to extend this idea to Slater basis set, Slater type orbitals (STO), instead of GTO orbitals, combined with LO or lo orbitals. This is done using shape functions.
@@ I Introduction
It is of paramount importance to make further progress on the basis set problem for Density Functional Theory (DFT) for small molecules. Small molecules larger then diatomic: carbon dioxide, water, ammonia, methane, ethyne, ethene, ethane to name a few have a variety of uses in industry sciences and engineering [2, 1]. There has been as much progress in optimizing basis sets for small molecule DFT calculations, similarly to the basis sets that have paved breakthroughs in solid state DFT [3, 4, 5, 9, 11, 12, 17, 14, 13, 8, 7, 6, 1, 10, 15, 16] - for crystalline solids. Much progress has been made using ideas based on Linear Combinations of Atomic Orbitals (LCAO) [19, 2, 18], where a good basis for the entire molecule consists of linear combinations of basis functions for the individual atoms. There are many candidate LCAO basis sets:
@@
  • Slater basis sets where Slater Type Orbitals (STOs) are used in the LCAO method. These are spherical harmonics multiplied by exponentially decaying radial wave functions. It is possible to have multiple wave functions per angular momentum channel - so called multi-zeta basis sets. STO’s have advantages over GTOs as they have the correct exponentially decaying behavior at infinity and the right Kato cusp at the origin [18, 2], however four center integrals are much harder - usually leading to shape function methods.
@@
  • Numerical basis sets where one solves the Khon Sham problem numerically and uses these wave functions as basis wave functions.
Furthermore plane waves, outside the LCAO concept, are often used to study periodically arranged arrays of molecules, with achievable distances between the molecules on the order of 10 $nm$, so that interactions between the molecules may be well neglected as a first approximation. The most common electronic structure method for this situation (plane wave basis sets for arrays of molecules) is the Projected Augmented Waves (PAW) [24, 25, 3, 4] method.
Here we focus on the LCAO method. Motivated by pseudopotential methods in DFT we propose a new (LCAO) hybrid basis set. Pseudopotentials, at their core, map a problem into another problem. The exact all electron wavefunction for the system is mapped onto a smooth (pseudized) wavefunction that may well be represented by a small number of plane waves, that is it is not the all electron wavefunction but simplified. Mappings are then made between the pseudized solution and the all electron wavefunction leading to a solution of the Khon Sham (KS) problem. Two of the pre-eminent examples of pseudopotentials are the Phillips Kleinman and Projected Augmented Waves (PAW) pseudopotentials. Here we show that these pseudopotential methods are all electron methods in disguise, where the all electron problem basis set includes smooth plane wave wave functions (with a low cutoff) and specialized wave functions for orbitals near atomic nuclei. Motivated by this observation, and the extensive numerical success of the PAW method [3, 4], we propose new basis sets for small molecules, where localized LO or lo [28, 5, 6, 1] basis wave functions are augmented with Gaussian wave functions - GTOs - (where Gaussians have been extensively used for the problem and various configuration integrals, including for center ones, for them are well known [18, 2, 3] thereby simplifying the calculation of the four center problem for Coulomb interactions) or STO’s for enhanced accuracy (but with shape functions). This construction is done in a manner similar to Phillips-Kleinman and PAW all electron basis sets when represented as all electron basis sets (see Section II). Multiple types of LO (lo) basis wave functions are possible. Thereby we propose that this new hybrid basis set is efficient for DFT and Hartree-Fock (HF) calculations of small molecules.
@@ II Many pseudopotential methods are all electron methods in disguise (motivation)
In this section we reformulate PAW and the Phillips-Kleinman method as all electron methods. We show the basis set of these all electron methods is composed of plane waves and special wave functions that are chosen to represent states near the core (near the nuclei). We show the two methods to be equivalent in their all electron formulation - depending on the exact choice of the localized wave functions near the nuclei (cores).
@@ II.1 Phillips-Kleinman as an all electron method in disguise
We now recall the Orthogonalized Plane Wave (OPW) wave functions used for the Phillips-Kleinman method. We know that [6, 3]:
@@
$-\sum_{i}\left\langle\phi_{i}\left(\mathbf{r}\right)\mid\frac{1}{\sqrt{V}}\exp\left(i\left(\mathbf{k}+\mathbf{K}\right)\cdot\mathbf{r}\right)\right\rangle\left|\phi_{i\mu}\left(\mathbf{r}\right)\right\rangle$ (1)
Here $\left|\phi_{i\mu}\left(\mathbf{r}\right)\right\rangle$ are localized wave functions relevant to the single site problem and $\mu$ labels the atoms. Here $V$ is the volume of the primitive lattice cell and $\mathbf{k}$ is a wavevector in the first Brillouin zone while $\mathbf{K}$ is a reciprocal lattice vector. Now the main claim of the Phillips-Kleinman method (greatly numerically supported [6, 3]) is that for a reasonable number of plane waves the exact Khon Sham (KS) wavefunction $\left|\psi_{n}\right\rangle$ can be written as
$\left|\psi_{n}\right\rangle\cong\sum_{\mathbf{K}}c_{\mathbf{k}}^{\mathbf{K}}\left|\phi_{OPW}\left(\mathbf{k}+\mathbf{K}\right)\right\rangle$ (2)
for some reasonable cutoffs $\mathbf{K}_{max}$ that means that
@@
$\subset Span\left\{\left\{\left|\frac{1}{\sqrt{V}}\exp\left(i\left(\mathbf{k}+\mathbf{K}\right)\cdot\mathbf{r}\right)\right\rangle\right\},\left\{\left|\phi_{i\mu}\left(\mathbf{r}\right)\right\rangle\right\}\right\}$ (3)
@@
$\left\{\left\{\left|\frac{1}{\sqrt{V}}\exp\left(i\left(\mathbf{k}+\mathbf{K}\right)\cdot\mathbf{r}\right)\right\rangle\right\},\left\{\left|\phi_{i\mu}\left(\mathbf{r}\right)\right\rangle\right\}\right\}$ (4)
As such Phillips-Kleinman is an all electron method in disguise, this will motivate our basis set constructions in Section III. We shall see below that PAW is very similar.
@@ II.2.1 PAW review
In the PAW formulation of the KS problem, we study the highly oscillatory eigenwavefunction of the KS problem near the nucleus of an atom using plane waves. In order to obtain a low cutoff (which can be as low as ~60 Rydberg in energy) and as such a tolerable basis set we must have that the wavefunction we represent using plane waves (the pseudized wavefunction) is very smooth. The key idea of PAW is then to introduce:
@@
where $\hat{\mathcal{T}}$ is some linear transformation (specified below) and $\left|\psi_{n}\right\rangle$ is the exact Khon Sham wave functions (solutions of the KS equations) while $\left|\tilde{\psi}_{n}\right\rangle$ are the pseudized wave functions we can approximate efficiently using plane waves at low cutoffs. Now we construct a set of augmentation spheres and we wish for $\hat{\mathcal{T}}$ to be the identity outside the augmentation spheres (centered around nuclei) , so we write:
@@
Here $\mu$ are the atomic augmentation spheres. Now we want the pseudized wave functions $\left|\tilde{\psi}_{n}\right\rangle$ to be smooth inside the sphere and we know that the exact KS wavefunction can be reasonably efficiently approximated by a single atom, atomic wavefunction $\left|\phi_{i\mu}\right\rangle$ near the atomic nuclei. As such if we choose $S_{\mathbf{\mu}}$ to be
$S_{\mu}\left|\tilde{\psi}_{n}\right\rangle=\sum c_{i\mu}^{n}\left|\phi_{i\mu}\right\rangle-\sum c_{i\mu}^{n}\left|\tilde{\phi}_{i\mu}\right\rangle$ (7)
where $\left|\tilde{\phi}_{i\mu}\right\rangle$ are smooth wave functions. Where we choose
$c_{i\mu}^{n}=\left\langle\tilde{p}_{i\mu}\mid\tilde{\psi}_{n}\right\rangle$ (8)
Here $\left|\tilde{p}_{i\mu}\right\rangle$ are the projector wave functions. As such we have that:
$\left|\psi_{n}\right\rangle=\hat{\mathcal{T}}\left|\tilde{\psi}_{n}\right\rangle=\left|\tilde{\psi}_{n}\right\rangle+\sum_{i}\left(\left|\phi_{i\mu}\right\rangle-\left|\tilde{\phi}_{i\mu}\right\rangle\right)\left\langle\tilde{p}_{i\mu}\mid\tilde{\psi}_{n}\right\rangle$ (9)
Furthermore the pseudized wave functions $\left|\tilde{\psi}_{n}\right\rangle$ may be well represented by plane waves.
@@ II.2.2 Reformulation of PAW as an all electron method
$\left|\psi_{n}\right\rangle\cong\left|\tilde{\psi}_{n}\right\rangle+\sum_{i}\left(\left|\phi_{i\mu}\right\rangle-\left|\tilde{\phi}_{i\mu}\right\rangle\right)\left\langle\tilde{p}_{i\mu}\mid\tilde{\psi}_{n}\right\rangle.$ (10)
@@
$\left|\psi_{n}\right\rangle\in Span\left\{\left\{\frac{1}{\sqrt{V}}\exp\left(i\left(\mathbf{k}+\mathbf{K}\right)\cdot\mathbf{r}\right)\right\},\left\{\left|\phi_{i\mu}\right\rangle\right\},\left\{\left|\tilde{\phi}_{i\mu}\right\rangle\right\}\right\}$ (11)
@@
$\left|\tilde{\phi}_{i\mu}\right\rangle\in Span\left\{\frac{1}{\sqrt{V}}\exp\left(i\left(\mathbf{k}+\mathbf{K}\right)\cdot\mathbf{r}\right)\right\}$ (12)
@@
$\left|\psi_{n}\right\rangle\in Span\left\{\left\{\frac{1}{\sqrt{V}}\exp\left(i\left(\mathbf{k}+\mathbf{K}\right)\cdot\mathbf{r}\right)\right\},\left\{\left|\phi_{i\mu}\left(\mathbf{r}\right)\right\rangle\right\}\right\}$ (13)
for a reasonable number of plane waves. Since the basis set of PAW and Phillips-Kleinman when formulated as all electron methods are similar, both methods should yield, in their all electron formulation, similar accuracy and require basis sets of the same size. They are equivalent within the all electron formulation when they use the same localized orbitals $\left|\phi_{i\mu}\left(\mathbf{r}\right)\right\rangle$.
@@ III GLO, SLO and PLO basis sets
Here we would like to work backwards, we wish to formulate an all electron method as the inverse of a pseudopotential method where we use specialized wave functions (with high wave numbers) near the nuclei and many regular, smooth, wave functions for the whole space. We focus on small molecules and crystalline solids. Below we will simply describe the final answer - that is relevant basis sets.
@@ III.1 Glo and GLO for small molecules
Here G in Glo and GLO stands for Gaussians while lo and LO for localized orbitals and the two types of elements (Gaussians and LO (lo) basis wave functions) are combined into one basis set. We consider the lo basis set [28]:
@@
where $S_{l}^{\mu}$ is chosen to make the wavefunction continuous everywhere [28]. Here:
@@
and $\bar{V}_{KS}\left(r\right)$ is the spherically average KS potential or HF single particle potential. Furthermore $E_{l}$ is chosen at the binding energy of the single atom system (other options are possible for example one could use the eigen-energies from the previous iteration to self consistency in the KS problem [26, 27] for the energies $E_{l}$). Alternatively we can use LO basis sets [28]:
@@
where $B_{l}^{\mu}$ and $S_{l}^{\mu}$ are chosen to make the wavefunction continuous and continuously differentiable everywhere [28]. Here $\dot{u}_{l}^{E_{l}}=\frac{\partial}{\partial E}u_{l}^{E_{l}}$. Higher derivative terms are also possible [28], also regular LO and lo is possible [5, 6, 1] as well as atom DFT based wave functions that is solutions to the radially averaged KS problem for a single atom [3]. Based on our experience with PAW a very small number of wave functions such as in Eq. (16) will suffice. We also add a large number of Gaussians to our basis set:
@@
to the molecular basis set. The wave functions $G_{l,m,\sigma}^{\mu}$, are smooth and replace the plane waves we have seen before in all electron formulations of pseudopotential methods. Here $\mathcal{N}_{l}$ is a normalization constant and cartesian Gaussians are also possible [20, 18]. This is a basis set for many small molecules. Furthermore Gaussians Coulomb integrals may be efficiently computed, see [30, 18] and Appendix A so there are “few” Coulomb integrals involving complicated functions [31]. Furthermore Eqs. (32) may be used, in the case when there is only one difficult to handle LO (lo) wavefunction, on two of the remaining Gaussians thereby simplifying that case to a three center problem. Overall if we consider the number of LO or lo orbitals as order one (highly reasonable for small molecules based on or experience with PAW [24, 25, 3]) this becomes a three center problem at worst. We note that contracted Gaussian Functions (CGF) are also possible [29, 18].